Integration Procedures(Itevrations by Parts)


∫ x³dx/cuberoot of (8-x²) from 0 to √7


1
Expert's answer
2021-04-26T17:40:33-0400


we evaluate the integral

"\int^{\sqrt7}_{0} \frac {x^3}{\sqrt[3]{8-x^2}}dx"

letting "u=8-x^3 ...(i)"

then "du= -2xdx" or"\frac {-1}{2}du=xdx ...(ii)"

for the limits; when "x=0, u= 8 ...(iii), 
x=\sqrt{7}, u=1 ...(iv)"

"\therefore \int^{\sqrt7}_{0} \frac{x^3}{\sqrt[3]{8-x^2}dx}=\int^{\sqrt7}_{0} \frac{x^2}{\sqrt[3]{8-x^2}}xdx"

"=\frac{-1}{2}\int^{1}_{8}\frac{(8-u)}{\sqrt[3]u}du" from (i) and (iv)

"=\frac{-1}{2}\int^{1}_{8} (8u^{-1/3}-u^{2/3})du"

"=\frac{-1}{2}[12u^{2/3}-\frac{3}{5}u^{5/3}]^{1}_{8}"

"=\frac {-1}{2} [\frac{57}{5}-\frac{144}{5}]=\frac{87}{10}=8.7"